# Experimental demonstration of the absence of noise-induced barren plateaus using information content landscape analysis

### v1

@@ Abstract
Variational quantum algorithms are a very promising tool for near-term quantum computing. However, despite their flexibility and wide applicability, their performance is fundamentally limited by Barren Plateaus (BP), where gradients vanish and optimization becomes intractable. Noise-Induced Barren Plateaus (NIBP) are particularly interesting, as they are predicted to arise due to noise accumulation independent of circuit structure. We experimentally study NIBP on IBM quantum hardware and demonstrate their absence under non-unital amplitude damping characterized by the qubit’s $T_{1}$ coherence times. We use Information Content Landscape Analysis (ICLA) to efficiently estimate gradient norms for circuits ranging from 8 to 102 qubits, with hundreds of parameters and circuit runtimes of hundreds of microseconds. Classical simulations of the 8-qubit case under noiseless, depolarizing, amplitude damping, and dephasing noise models serve as a baseline comparison. We thoroughly analyze the experimental results considering calibration data, shot-noise, and circuit structure. We robustly observe that the gradient magnitude saturates beyond a characteristic circuit runtime, in contrast with the exponential decay expected from NIBP. Using recent theoretical results, we corroborate that under $T_{1}$-dominated noise NIBP do not occur and extract an effective $T^{{\rm eff}}_{1}$ that is significantly shorter than suggested by standard calibration data. Our results experimentally confirm recent predictions on the absence of NIBP under non-unital noise. These findings also indicate that conventional benchmarking metrics based on average values for device characteristics may be insufficient to predict variational algorithm performance, but full distributions need to be considered.
@@ I Introduction
The manuscript is organized as follows: In Section˜II, we provide the necessary background for non-expert readers which include a description of variational quantum algorithms Section˜II.1, information content landscape analysis Section˜II.2, a basic introduction to noise in quantum hardware and noise-induced barren plateaus Section˜II.3, and finally a general theory of barren plateaus Section˜II.4. An expert reader might skip some or all of these subsections. Next, in Section˜III we detail the experimental set-up, and the parameters of the IBM quantum processing units. The main results of our work are presented in Section˜IV. Finally, in Section˜V we discuss the implications of this study, lay down open research questions, and present some research directions for the future.
II Background
@@ II.4 Barren Plateaus
III Experimental setup on IBM quantum hardware
@@
We transpile logical circuits given by Equation˜15 with default setting on the IBM qiskit transpiler. Detail on the gate statistics and circuit depths, as well as representative examples of the circuit used in this study are given in Appendix˜A.
We used ICLA to estimate the norm of the gradient of the cost landscape for $N\in\{8,20,45,65,102\}$ qubits and number of layers up to $L=120$. The landscape was created using $M=\min(10m,200)$ parameter vectors, where $m=2L$ is determined by the number of layers of the quantum circuits. We truncated the size of the landscape to $M=200$ for large systems as we observed this to be sufficient. See Appendix˜B for more details on this. This observation makes the ICLA even more efficient as it substantially reduces the number of required circuit evaluations for large circuit depth $L$.
We estimate the cost function value by using $R=16384$ measurement shots for each parameter set. We ensured that the shot-noise error was always at least a factor of two below the observed gradient signal. This forced us to perform some few experiments with $R=32768$ and $R=65536$ measurement shots. We refer the reader to Appendix˜C for more details on shot-noise floor and how cost function values are calculated.
@@
We ran experiments on IBM $127$-qubit Falcon processors ibm_brisbane, ibm_kyiv, and ibm_sherbrooke, as well as the 156 qubit Heron processor ibm_fez. For each experimental run we recorded the assignment of logical to physical qubits and the backend calibration data, in particular $T_{1}$, $T_{2}$, ECR-gate error, and readout error. More details on the hardware calibration data can be found in Appendix A, in particular Table. 2.
IV Results
IV.1 Numerical and experimental evidence of the absence of NIBP
@@
We start by discussing the results of the classical simulations (dashed lines). For the simulated circuits we estimate the hardware runtime $t_{cir}$ by taking their depth and translating it using the average time needed for a given circuit depth from the hardware runs (see Appendix˜A for more details). The noiseless simulation (dark purple) does not show any sign of BP indicated by $\norm{\nabla C}/C_{0}$ initially decreasing from $0.5$, but then leveling off at $0.05$ for runtimes above $\sim 150\mu s$. This is the expected behavior in our experiment based on our choice of a local observable to avoid cost-function dependent BP [17]. Conversely, NIBP arise in the noisy simulations employing a single-qubit depolarizing noise model of Eq. (9). A clear exponential decay in $\norm{\nabla C}/C_{0}$ going from $1$ to $10^{-4}$ is observed in a very short runtime, $t_{cir}\lesssim 150\,\mu s$, for single-qubit depolarizing probability, $p=0.025$ (light green). The effect is less severe for a milder depolarizing probability, $p=0.005$, where it takes approximately twice as long to reach the same value of $\norm{\nabla C}/C_{0}$ (dark green). In any case, we conclude that the depolarizing noise model inevitably leads to NIBP as extensively analyzed in [16]. We also simulate the experiment under a more realistic noise model that includes $T_{1}$ and $T_{2}$ coherence times as described in Eq. (12) (light blue). We used randomized coherence times drawn from normal distributions $T_{1}\sim\mathcal{N}(\mu=244\mu s,\sigma=74\mu s)$ and $T_{2}\sim\mathcal{N}(\mu=159\mu s,\sigma=93\mu s)$ for this noise model, which are consistent with the coherence times from the hardware experiments. At short circuit times, up to $t_{cir}\lesssim 170\,\mu s$, both the realistic noise and the weak depolarizing noise models have an almost identical decay, starting at $\norm{\nabla C}/C_{0}\approx 0.5$ and exponentially decreasing to $\norm{\nabla C}/C_{0}\approx 0.005$. At $t_{cir}\gtrsim 170\,\mu s$ to $t_{cir}\approx 230\,\mu s$ the trend of $\norm{\nabla C}/C_{0}$ in the realistic noise model changes drastically, and at $t_{cir}\gtrsim 230\,\mu s$ onwards, it fully flattens to $\norm{\nabla C}/C_{0}\approx 0.002$. These simulations confirm that non-unital noise in the form of amplitude damping prevents the appearance of NIBP for runtimes on the order of the decoherence time, $t_{cir}\sim T_{1}$.
@@
In order to give some physical intuition for the vanishing NIBP, a spectral analysis of the simulated 8-qubit density matrices Appendix˜D was carried out. The spectrum of the depolarizing noise model indicated a fully mixed state for large circuit depth (long runtimes). On the other hand, the spectrum of the amplitude damping noise model was confined to a noise-induced limit set as described in ref. [20]. The noise-induced limit set is characterized by a density matrix where the spectral weight is distributed over a finite subset of states. Because the noise-induced limit set depends on the variational parameters of the quantum circuit, the corresponding cost function values also depend on the variational parameters. Therefore, the noise-induced limit set leads to cost function values which are not exponentially concentrated around one value, therefore, a finite gradient as we observe.
@@
IV.2 Effective $T_{1}$ from the absence of NIBP
@@ 65 $247\pm 77$ 56 $220\pm 15$ $66\pm 51 11$
We validated these insights by running more experiments with different setups. First, we ran the same circuits on the different hardware platforms. The ibm_kyiv and ibm_sherbrooke platform are the same Falcon architecture as ibm_brisbane and only differ in the distribution of coherence times and errors. In contrast, the ibm_fez platform is of the Heron architecture where different gates with much shorter gate times are implemented. However, the coherence times of ibm_fez are also substantially shorter as can be seen in Appendix˜A. Second, we ran the circuits more efficient gate sequence as described in Appendix˜A, which leads to much shorter runtime on ibm_brisbane for $N=45$. And lastly, we also implemented a slightly modified Hamiltonian where the coupling topology was exactly that of the heavy hex hardware coupling.
Figure 3: Gradient, $\norm{\nabla C}/C_{0}$, as a function of circuit times $t_{cir}$ for various setups. The data labeled ’short’ refers to the different compilation of the same logical circuit (see Appendix˜A), while ’heavy hex’ indicates that a different Hamiltonian with same coupling topology as the hardware was used. These results show that the flattening phenomenon is consistent across different hardware and setups.
@@ V Conclusions
However, the mere fact the a finite gradient is retained under amplitude damping, does not imply that such circuits are useful for solving a given task. As remarked above and in ref. [21], non-unital noise implies a finite and shallow effective circuit depth, which limits the expressive power, and therefore does not allow for the solution of complex problems. In addition, only the variational parameters of the later layers determine the cost function values, and the influence of the early layers is lost. This calls for re-evaluating insights related to parameter concentration and parameter transfer in variational quantum circuits [44, 45, 46, 41, 42], where warm-starting schemes based on transferring parameters or optimization-free schemes might not lead to increased success probabilities. Yet, recent work suggest that non-unital noise might be useful in reservoir computing [47].
It is still an interesting research topic how errors from non-unital noise channels like amplitude damping can be mitigated on near-term hardware, since popular schemes like dynamical decoupling, Pauli twirling and zero-noise extrapolation are not effective. Error mitigation by symmetry verification [48] is efficient as it can be done in postprocessing, but it is restricted to specific problems. Approaches based on weak measurements [49, 50, 51] also seems promising and might be easier to implement than full error correction schemes [52, 53, 54].
We leave all these interesting research directions for future work.
Acknowledgments
The authors would like to thank Phillip Hauke, Adrián Pérez-Salinas, and Stefano Polla for stimulating and fruitful discussions. A.B. acknowledges funding from the Honda Research Institute Europe GmbH, Provincia Autonoma di Trento and Q@TN, the joint laboratory between University of Trento, FBK—Fondazione Bruno Kessler, INFN—National Institute for Nuclear Physics, and CNR—National Research Council. S.S. and L.E. acknowledge funding by the European Union under Horizon Europe Programme, Grant Agreement 101080086–NeQST.
Author contribution
Appendix A Details on hardware experiments
A.1 Circuit structure
Figure 5: Examples of the same logical $L=1$ single layer circuit with $N=20$ qubits transpiled in the two different ways. Ladder structure (left) which is the default structure we used and parallel execution (right) which we termed ’short’.
@@ A.1 Circuit structure
A.2 Circuit depth and runtimes
@@ Appendix B Robustness of ICLA
Appendix C Experimental details
@@
C.1 Estimating cost function values
Figure 9: Sorted cost values for $N=20$ (upper) and $N=102$ (lower) from the experiments of the ibm_brisbane device. The error bars indicate the variation expected for to shot-noise with $R=16384$ shots used here. The insets show cost spectra for parameter values producing the minimal and maximal cost function value.
@@ C.2 Noise influences
Appendix D Density matrix analysis
@@
Although limited to small system sizes, these simulations highlight how unital (depolarizing) noise drives the system towards the maximally mixed state and induces barren plateaus, while non unital (amplitude damping) noise preserves spectral structure and leads to a noise-induced limit set with finite average gradients.

### v2

@@ Abstract
Variational quantum algorithms are promising candidates for near-term quantum computing but can be hindered by barren plateaus, where gradients vanish exponentially and optimization becomes intractable. Noise-Induced Barren Plateaus (NIBP) are particularly concerning because they are predicted to arise generically from noise accumulation, independent of system size, circuit structure, and observable locality. We experimentally investigate NIBP on IBM quantum hardware. Using Information Content Landscape Analysis (ICLA), we efficiently estimate gradient norms for variational circuits ranging from 8 to 102 qubits, up to hundreds of parameters and circuit runtimes of hundreds of microseconds. Contrary to NIBP expectations, we observe that gradient magnitudes saturate beyond a characteristic runtime rather than decaying exponentially. Classical simulations of the 8-qubit case under noiseless, depolarizing, amplitude-damping, and dephasing noise models support this behavior. Consistent with recent theory, our results show that $T_{1}$-dominated non-unital noise inhibits the emergence of NIBP. Our analysis suggest that average calibration metrics may be insufficient to predict variational algorithm performance.
@@ I Introduction
The manuscript is organized as follows: In the first part of Section˜II, we provide the necessary background for non-expert readers which include a description of variational quantum algorithms Section˜II.1, information content landscape analysis Section˜II.2, a basic introduction to noise in quantum hardware and noise-induced barren plateaus Section˜II.3, and finally a general theory of barren plateaus Section˜II.4. An expert reader might skip some or all of these subsections. The second half of Section˜II is devoted to the experiment and the main results of this work. In Section˜II.5 we detail the experimental set-up, and the parameters of the IBM quantum processing units, followed by the presentation of the main figures in Section˜II.6, and the connection of the abscense of NIBP to an effective coherence time in Section˜II.7. The manuscript continues with Section˜III where we discuss the implications of this study, lay down open research questions, and present some research directions for the future. Finally, in Section˜IV we present further details of the experimental set up to ease its reproducibility, as well as additional calculation to support our main results.
II Results
@@ II.4 Barren Plateaus
II.5 Experimental setup on IBM quantum hardware
@@
We transpile logical circuits given by Equation˜15 with default setting on the IBM qiskit transpiler. Detail on the gate statistics and circuit depths, as well as representative examples of the circuit used in this study are given in Section˜IV.1.
We used ICLA to estimate the norm of the gradient of the cost landscape for $N\in\{8,20,45,65,102\}$ qubits and number of layers up to $L=120$. The landscape was created using $M=\min(10m,200)$ parameter vectors, where $m=2L$ is determined by the number of layers of the quantum circuits. We truncated the size of the landscape to $M=200$ for large systems as we observed this to be sufficient. See Section˜IV.4 for more details on this. This observation makes the ICLA even more efficient as it substantially reduces the number of required circuit evaluations for large circuit depth $L$.
We estimate the cost function value by using $R=16384$ measurement shots for each parameter set. We ensured that the shot-noise error was always at least a factor of two below the observed gradient signal. This forced us to perform some few experiments with $R=32768$ and $R=65536$ measurement shots. We refer the reader to Section˜IV.5 for more details on shot-noise floor and how cost function values are calculated.
@@
We ran experiments on IBM $127$-qubit Falcon processors ibm_brisbane, ibm_kyiv, and ibm_sherbrooke, as well as the 156 qubit Heron processor ibm_fez. For each experimental run we recorded the assignment of logical to physical qubits and the backend calibration data, in particular $T_{1}$, $T_{2}$, ECR-gate error, and readout error. More details on the hardware calibration data can be found in Appendix IV.1, in particular Table. 2.
II.6 Numerical and experimental evidence of the absence of NIBP
@@
We start by discussing the results of the classical simulations (dashed lines). For the simulated circuits we estimate the hardware runtime $t_{cir}$ by taking their depth and translating it using the average time needed for a given circuit depth from the hardware runs (see Section˜IV.1 for more details). The noiseless simulation (dark purple) does not show any sign of BP indicated by $\norm{\nabla C}/C_{0}$ initially decreasing from $0.5$, but then leveling off at $0.05$ for runtimes above $\sim 150\mu s$. This is the expected behavior in our experiment based on our choice of a local observable to avoid cost-function dependent BP [17]. Conversely, NIBP arise in the noisy simulations employing a single-qubit depolarizing noise model of Eq. (9). A clear exponential decay in $\norm{\nabla C}/C_{0}$ going from $1$ to $10^{-4}$ is observed in a very short runtime, $t_{cir}\lesssim 150\,\mu s$, for single-qubit depolarizing probability, $p=0.025$ (light green). The effect is less severe for a milder depolarizing probability, $p=0.005$, where it takes approximately twice as long to reach the same value of $\norm{\nabla C}/C_{0}$ (dark green). In any case, we conclude that the depolarizing noise model inevitably leads to NIBP as extensively analyzed in [16]. We also simulate the experiment under a more realistic noise model that includes $T_{1}$ and $T_{2}$ coherence times as described in Eq. (12) (light blue). We used randomized coherence times drawn from normal distributions $T_{1}\sim\mathcal{N}(\mu=244\mu s,\sigma=74\mu s)$ and $T_{2}\sim\mathcal{N}(\mu=159\mu s,\sigma=93\mu s)$ for this noise model, which are consistent with the coherence times from the hardware experiments. At short circuit times, up to $t_{cir}\lesssim 170\,\mu s$, both the realistic noise and the weak depolarizing noise models have an almost identical decay, starting at $\norm{\nabla C}/C_{0}\approx 0.5$ and exponentially decreasing to $\norm{\nabla C}/C_{0}\approx 0.005$. At $t_{cir}\gtrsim 170\,\mu s$ to $t_{cir}\approx 230\,\mu s$ the trend of $\norm{\nabla C}/C_{0}$ in the realistic noise model changes drastically, and at $t_{cir}\gtrsim 230\,\mu s$ onwards, it fully flattens to $\norm{\nabla C}/C_{0}\approx 0.002$. These simulations confirm that non-unital noise in the form of amplitude damping prevents the appearance of NIBP for runtimes on the order of the decoherence time, $t_{cir}\sim T_{1}$.
@@
In order to give some physical intuition for the vanishing NIBP, a spectral analysis of the simulated 8-qubit density matrices Section˜IV.7 was carried out. The spectrum of the depolarizing noise model indicated a fully mixed state for large circuit depth (long runtimes). On the other hand, the spectrum of the amplitude damping noise model was confined to a noise-induced limit set as described in ref. [20]. The noise-induced limit set is characterized by a density matrix where the spectral weight is distributed over a finite subset of states. Because the noise-induced limit set depends on the variational parameters of the quantum circuit, the corresponding cost function values also depend on the variational parameters. Therefore, the noise-induced limit set leads to cost function values which are not exponentially concentrated around one value, therefore, a finite gradient as we observe.
@@
II.7 Effective $T_{1}$ from the absence of NIBP
@@ 65 $247\pm 77$ 56 $220\pm 15$ $66\pm 51 11$
We validated these insights by running more experiments with different setups. First, we ran the same circuits on the different hardware platforms. The ibm_kyiv and ibm_sherbrooke platform are the same Falcon architecture as ibm_brisbane and only differ in the distribution of coherence times and errors. In contrast, the ibm_fez platform is of the Heron architecture where different gates with much shorter gate times are implemented. However, the coherence times of ibm_fez are also substantially shorter as can be seen in Section˜IV.1. Second, we ran the circuits more efficient gate sequence as described in Section˜IV.1, which leads to much shorter runtime on ibm_brisbane for $N=45$. And lastly, we also implemented a slightly modified Hamiltonian where the coupling topology was exactly that of the heavy hex hardware coupling.
Figure 3: Gradient, $\norm{\nabla C}/C_{0}$, as a function of circuit times $t_{cir}$ for various setups. The data labeled ’short’ refers to the different compilation of the same logical circuit (see Section˜IV.1), while ’heavy hex’ indicates that a different Hamiltonian with same coupling topology as the hardware was used. These results show that the flattening phenomenon is consistent across different hardware and setups.
@@ III Discussion
However, the mere fact the a finite gradient is retained under amplitude damping, does not imply that such circuits are useful for solving a given task. As remarked above and in ref. [21], non-unital noise implies a finite and shallow effective circuit depth, which limits the expressive power, and therefore does not allow for the solution of complex problems. In addition, only the variational parameters of the later layers determine the cost function values, and the influence of the early layers is lost. This calls for re-evaluating insights related to parameter concentration and parameter transfer in variational quantum circuits [44, 45, 46, 41, 42], where warm-starting schemes based on transferring parameters or optimization-free schemes might not lead to increased success probabilities. Yet, recent work suggest that non-unital noise and dissipative processes can be leveraged for useful computation [47] as well as in reservoir computing [48].
It is still an interesting research topic how errors from non-unital noise channels like amplitude damping can be mitigated on near-term hardware, since popular schemes like dynamical decoupling, Pauli twirling and zero-noise extrapolation are not effective. Error mitigation by symmetry verification [49] is efficient as it can be done in postprocessing, but it is restricted to specific problems. Approaches based on weak measurements [50, 51, 52] also seem promising and might be easier to implement than full error correction schemes [53, 54, 55]. All of these are interesting research directions that we leave for future work.
IV Methods
In this section we provide details necessary reproduce the numerical and experimetnal results in the mansucript.
IV.1 Circuit structure
Figure 5: Examples of the same logical $L=1$ single-layer circuit with $N=20$ qubits with two different transpilations. Ladder structure (left) which is the default structure used in our experiment and parallel execution (right) which we termed ’short’.
@@ IV.1 Circuit structure
IV.2 Circuit depth and runtimes
@@ IV.4 Robustness of ICLA
IV.5 Estimating cost function values
@@
Figure 9: Sorted cost values for $N=20$ (upper) and $N=102$ (lower) from the experiments of the ibm_brisbane device. The error bars indicate the variation expected for to shot-noise with $R=16384$ shots used here. The insets show cost spectra for parameter values producing the minimal (orange) and maximal (green) cost function values for small (upper) and large (lower) number of layers. The blue curve is the spectrum of the Hamiltonian.
@@ IV.6 Noise influences
IV.7 Density matrix analysis
@@
Although limited to a small system size, these simulations highlight how unital (depolarizing) noise drives the system towards the maximally mixed state and induces barren plateaus, while non unital (amplitude damping) noise preserves spectral structure and leads to a noise-induced limit set, which leads to finite average gradients.
Acknowledgments
The authors would like to thank Phillip Hauke, Adrián Pérez-Salinas, Stefano Polla, Daniel Lidar, and Phattharaporn Singkanipa for stimulating and fruitful discussions. The authors also would like to thank Bernhard Sendhoff, Jens Schmuedderich, Heiko Wersing, Avetik Harutyunyan, and the Honda Research Institute network for support in this project. A.B. acknowledges funding from the Honda Research Institute Europe GmbH, Provincia Autonoma di Trento and Q@TN, the joint laboratory between University of Trento, FBK—Fondazione Bruno Kessler, INFN—National Institute for Nuclear Physics, and CNR—National Research Council. S.S. and L.E. acknowledge funding by the European Union under Horizon Europe Programme, Grant Agreement 101080086–NeQST.
Author contribution
Competing interests
The authors declare no competing interests.
Data availability
The data supporting the findings of this study are available from the corresponding author upon reasonable request.
Code availability
All packages used in this work are open source and available via git repositories online.